Dy/dx+{x/(1-x^2)}y=x√y,y(0)=1
1
Expert's answer
2020-04-20T11:56:11-0400

"\frac {dy}{dx}+\frac {x}{1-x^2}y=x\sqrt{y} , y(0)=1"

Trivial solution "y=0", but there is "y(0)=1" . That's why we can't take it.

Further we consider that "y\neq0."

"y'+\frac{x}{1-x^2}y=x\sqrt{y}\\
\frac{y'}{\sqrt{y}}+\frac{x\sqrt{y}}{1-x^2}=x"


Let "\sqrt{y}=z", then "z'=(\sqrt{y})'=\frac{y'}{2\sqrt{y}}" and "\frac{y'}{\sqrt{y}}=2z'"

"2z'+\frac{xz}{1-x^2}=x"


Let "z=uv", then "z'=u'v+uv'"

"2u'v+2uv'+\frac{xuv}{1-x^2}=x\\
2u'v+u(2v'+\frac{xv}{1-x^2})=x" (1)


"2v'+\frac{xv}{1-x^2}=0\\
2v'=\frac{xv}{x^2-1}\\
2\frac{dv}{dx}=\frac{xv}{x^2-1}\\
2\frac{dv}{v}=\frac{xdx}{x^2-1}\\
2\int\frac{dv}{v}=\int\frac{xdx}{x^2-1}\\
2log(|v|)=1/2log(|x^2-1|)\\
v=\sqrt[4]{|x^2-1|}"


Now (1) is "2u'\sqrt[4]{|x^2-1|}=x"

"2\frac{du}{dv}=\frac{x}{\sqrt[4]{|x^2-1|}}\\
2du=\frac{xdx}{\sqrt[4]{|x^2-1|}}\\
2\int du=\int \frac{xdx}{\sqrt[4]{|x^2-1|}}\\
2\int du=\frac{1}{2}\int \frac{dx^2}{\sqrt[4]{|x^2-1|}}\\
u=\frac{1}{3}\sqrt[4]{|x^2-1|}^3+C"


Finally

"z=uv=\sqrt[4]{|x^2-1|}\frac{1}{3}(\sqrt[4]{|x^2-1|}^3+C)=\frac{1}{3}(|x^2-1|)+C\sqrt[4]{|x^2-1|}"

"\sqrt{y}=\frac{1}{3}(|x^2-1|)+C\sqrt[4]{|x^2-1|}\\
y=(\frac{1}{3}(|x^2-1|)+C\sqrt[4]{|x^2-1|})^2"


"y(0)=1, \implies \\
1=(\frac{1}{3}|-1|+C\sqrt[4]{|-1|})^2\\
1=(\frac{1}{3}+C)^2"

"C=\frac{2}{3}" or "C=-\frac{4}{3}"


"y=(\frac{1}{3}(|x^2-1|)+\frac{2}{3}\sqrt[4]{|x^2-1|})^2" or

"y=(\frac{1}{3}(|x^2-1|)-\frac{4}{3}\sqrt[4]{|x^2-1|})^2"